OpenStudy (mythorius):

A right triangle has an acute angle measure of 22°. Which two numbers represent the lengths of the legs of this triangle? 1 and 5 2 and 5 3 and 5 4 and 5 @KamiBug

OpenStudy (mythorius):

@ILovePuppiesLol

zepdrix (zepdrix):

\[\large\rm \tan22^o\approx 0.4=\frac{4}{10}\]

zepdrix (zepdrix):

So which tangent ratio would get you that value?\[\large\rm \tan22^o=\frac{opposite}{adjacent}\]

zepdrix (zepdrix):